# Non-transitive pseudo-Anosov flows

### v1

@@ 3. Smale classes and topological transitivity
As above, consider $\alpha$, $\beta$ fixing all half leaves through $a$ and $b$ that are expanding on $\mathcal{F}^{+}(a)$ and $\mathcal{F}^{+}(b)$, respectively; and hence preserving local orientation in the rectangle bounded by ±$\mathcal{F}^{\pm}(a)$ and ±$\mathcal{F}^{\pm}(b)$. Proposition 3.9, taking $g$ to be the identity, shows that for $n,m$ sufficiently large $\alpha^{n}\beta^{m}$ has a fixed point $z$ in any given neighborhood $U$ of $\mathcal{F}^{+}(a)\cap\mathcal{F}^{-}(b)$, and the proof (considering the images of $I$ and $J$) easily shows that this necessarily lies in the quadrilateral bounded by ±$\mathcal{F}^{\pm}(a)$ and $\mathcal{F}^{\mp}(b)$. Similarly, given a neighborhood $U^{\prime}$ of $\mathcal{F}^{+}(b)\cap\mathcal{F}^{-}(a)$ in the quadrilateral bounded by the leaves of $a,b$, for sufficiently large $m,n$ the map $\beta^{m}\alpha^{n}$ has a fixed point $z^{\prime}$ in $U^{\prime}$. Since $\beta^{m}\alpha^{n}=\beta^{m}(\alpha^{n}\beta^{m})\beta^{-m}$, we have that $z^{\prime}=\beta^{n}(z)$, as desired. ∎
@@
One easily shows that, up to shrinking the product neighborhood in the above corollary, the closed sets $C_{i}$ are either a closed interval, a Cantor set or an isolated point. The work in the next few sections will show that, as in the case of Anosov flows, only four of the six (up to symmetry) possible topological types of product can happen: One may have a product $\mathrm{Interval}\times\mathrm{Interval}$ (which happens only if the action is transitive), a product $\mathrm{Interval}\times\mathrm{Cantor}$ (which happens in extremal Smale classes, see Observation 6.3), a product $\mathrm{Cantor}\times\mathrm{Cantor}$ (which happens for any non-isolated point in a non-extremal Smale class), and finally isolated fixed points, i.e., of the form $\mathrm{Isolated}\times\mathrm{Isolated}$. See Proposition 4.12 and Corollary 4.14 for a proof.
@@
Using this, we can prove Theorem 1.8. Recall the statement:
See 1.8
@@
We will show that for any open sets $U,V$ such that $U\cap\Lambda\neq\emptyset$ and $V\cap\Lambda\neq\emptyset$, there exists $g\in G$ such that $gU\cap V\cap\Lambda\neq\emptyset$. This implies the existence of a dense orbit (the classical proof for continuous maps applies here, see e.g. [KH95, Lemma 1.4.2], since Smale classes are closed in the complement of the singular leaves in $P$).
@@
In the special case where there is a unique Smale class $\Lambda$, Axiom (A2) together with Proposition 3.8 implies that $\bar{\Lambda}=P$, so $G$ is topologically transitive on $P$ and $P$ contains a dense subset of points fixed by nontrivial elements, i.e., Axioms (A2a) and (A2b) are both satisfied. Conversely, if $\mathrm{Fix}_{G}$ is dense, then $\mathcal{R}_{G}$ is dense (since $\mathcal{R}_{G}=\mathrm{Fix}_{G}\smallsetminus\mathrm{Sing}$, and $\mathrm{Sing}$ is a closed discrete subset of $P$ by Lemma 2.21), moreover any product neighborhood intersects at most one Smale class, so there is a unique Smale class by connectedness of $P\smallsetminus\mathrm{Sing}$. ∎
Theorem 1.7 is not an immediate consequence.
@@ 4. Smale chains and the structure of wandering sets
For Anosov flows on compact manifolds, Smale classes corresponds to basic sets. By Brunella [Bru93] these sets are separated by a collection of tori transverse to the flow. Moreover, in the nontransitive case any basic set contains “boundary” orbits, which are periodic orbits such that (at least) one of their rays are contained in the wandering set (these are the free separatrices appearing in [BB02, Lemme 1.6]). A related theory is developed for Smale diffeomorphisms of surfaces in [BL98].
@@
Note that, by comparison, the first step of the standard proof of the existence of separating transverse tori is the existence of a Lyapunov function for Anosov flows. Furthermore, as mentioned in the introduction, in the case of pseudo-Anosov flows, here we detect more of the tori in the non-wandering set than what the proof using Lyapunov functions would, since Lyapunov functions only allow to separate chain transitive components, and not the components of the non-wandering set. See Proposition 8.8.
@@ 4.1. Structure of the regular and “wandering” sets
Equivalently, if $Q_{0},\dots,Q_{2p-1}$ are the $TL$ subquadrants, cyclically ordered, and $Q_{0}$ is wandering, then there exists $i$, $10\leq i\leq 2p-$ 1, such that both $Q_{i}$ and $Q_{i+1}$ are wandering.
@@
Up to renaming the $TL$ subquadrants, we assume that $Q_{0},\dots,Q_{2p-1}$ are ordered so that $Q_{0}$ shares a $\mathcal{F}^{+}$-side with $Q_{1}$ and a $\mathcal{F}^{-}$-side with $Q_{2p-1}$. If the $Q_{2i-1}$,$i=1,\dots,p$ are all regular, then there exist points $y_{i}\in\mathrm{Fix}(G)\cap Q_{2i-1}$. As in the proof of Lemma 4.5, after applying a high positive or negative power of the element of $G$ fixing $c$ (which exists by axiom (A3)), we can replace $y_{i}$ with points $y_{i}^{\prime}$ with ±$\mathcal{F}^{\pm}{y_{i}^{\prime}}\cap\mathcal{F}^{\mp}(c)$ as close as we like to $c$. By then applying Corollary 3.10 using $y_{i}^{\prime}$ and $c$, we can obtain points $x_{i}\in\mathrm{Fix}(G)\cap Q_{2i-1}$ as close as we like to $c$. In particular, we can take these sufficiently close to $c$ so that $\mathcal{F}^{+}(x_{i})\cap\mathcal{F}^{-}(x_{i+1})\neq\emptyset$ and $\mathcal{F}^{+}(x_{p})\cap\mathcal{F}^{-}(x_{1})\neq\emptyset$. Let $h_{i}\neq id$ be an element of $G$ that fixes $x_{i}$. Then for appropriate powers $n_{i}$, we have that $h_{p-1}^{n_{p-1}}\dots h_{2}^{n_{2}}\mathcal{F}^{+}(x_{1})\cap\mathcal{F}^{-}(x_{p})\neq\emptyset$. That is, we are in the setting of Proposition 3.8, and therefore there is a point fixed by some nontrivial element of $G$ arbitrarily close to $\mathcal{F}^{+}(x_{p})\cap\mathcal{F}^{-}(x_{1})\in Q_{0}$, contradicting the fact that $Q_{0}$ is wandering. ∎
@@
Suppose $x\in\overline{\mathrm{Fix}}_{G}$, and suppose that it has a wandering TL-substrip $S$ based on a ray $r_{0}$ of a leaf of $\mathcal{F}^{+}$. Then $r_{0}$ makes a perfect fit with a leaf of $\mathcal{F}^{-}$ on the side of the wandering substrip.
@@
Thanks to Lemma 4.8, we may assume that the boundary of $S$ consists of $r_{0}$, a ray $r^{\prime}_{1}$ of a leaf fixed by some element $g\in G$, a compact interval of a leaf of $\mathcal{F}^{-}$ between the two initial points of the rays, and a part of a leaf $l^{-}\in\mathcal{F}^{-}$. Moreover, Lemma 4.8 gives three possibilities for $l^{-}$. One of the three options is that $l^{-}$ makes a perfect fit with $r^{+}$, which is what we want to show. Hence, we have to eliminate the other two.
@@ 4.2. Boundary leaves
A boundary leaf is a leaf $l\in\mathcal{F}^{\pm}$ such that there exists $a\in l\cap\overline{\mathrm{Fix}}_{G}$ and $a$ is a boundary point of $\overline{\mathrm{Fix}}_{G}\cap\mathcal{F}^{\mp}(a)$. (Analogously, one can also define boundary faces for singular leaves.)
@@
Let $x\in\mathcal{R}_{G}$ be such that ±$\mathcal{F}^{\pm}(x)$ is an isolated boundary leaf, i.e., for some small enough product neighborhood $U$ of $x$, we have $\overline{\mathrm{Fix}}_{G}\cap U\subset\mathcal{F}^{\pm}(x)$. Then $x\in\mathrm{Fix}_{G}$ and it is the corner of four wandering lozenges. In particular, the Smale class of $x$ consists only of $x$ and its images under $G$.
@@
Let $U$ be a product neighborhood of $x$ such that $\overline{\mathrm{Fix}}_{G}\cap U\subset\mathcal{F}^{+}(x)$. (The proof for $\mathcal{F}^{-}(x)$ being isolated is obviously symmetric). Since $x\in\mathcal{R}_{G}$, it is accumulated by elements of $\mathrm{Fix}_{G}$, but any such element in $U$ must be in $\mathcal{F}^{+}(x)$ by assumption. Since any leaf contains at most one element of $\mathrm{Fix}_{G}$ (see Remark 1.3), we deduce that $x\in\mathrm{Fix}_{G}$. Hence $\overline{\mathrm{Fix}}_{G}\cap U=\{x\}$, so (by the product structure of $\overline{\mathrm{Fix}}_{G}$) every TL subquadrant of $x$ are wandering and the conclusion follows from Corollary 4.10. ∎
@@
Assuming $y=\infty$, by the above, both TL subquadrants of $x$ along the ray $(x,y)$ are wandering. In particular, the ray $(x,y)$ is the base of wandering TL-substrips on each sides. By Lemma 4.9 there exist leaves $l_{1},l_{2}$ of $\mathcal{F}^{-}$ making perfect fits on each side. These are nonseparated in the leaf space of $\mathcal{F}^{+}$, so are fixed by some nontrivial element $g\in G$ (Prop. 2.10), forming the sides of adjacent $g$-invariant lozenges (Corollary 2.19), with common shared side on $\mathcal{F}^{+}(x)$. Moreover, these lozenges are wandering, so the common corner on $\mathcal{F}^{+}(x)$ is the element in $\overline{\mathrm{Fix}}_{G}\cap\mathcal{F}^{+}(x)$ closest to the perfect fit, i.e., it is $x$. ∎
@@
    $I$ is unbounded, $a$ is fixed by $g$, and $a$ is the corner of two wandering lozenges with $I$ as their shared sides,

### v2

@@ 3. Smale classes and topological transitivity
As above, consider $\alpha$, $\beta$ fixing all half leaves through $a$ and $b$ that are expanding on $\mathcal{F}^{+}(a)$ and $\mathcal{F}^{+}(b)$, respectively; and hence preserving local orientation in the rectangle bounded by $\mathcal{F}^{\pm}(a)$ and $\mathcal{F}^{\pm}(b)$. Proposition 3.10, taking $g$ to be the identity, shows that for $n,m$ sufficiently large $\alpha^{n}\beta^{m}$ has a fixed point $z$ in any given neighborhood $U$ of $\mathcal{F}^{+}(a)\cap\mathcal{F}^{-}(b)$, and the proof (considering the images of $I$ and $J$) easily shows that this necessarily lies in the quadrilateral bounded by $\mathcal{F}^{\pm}(a)$ and $\mathcal{F}^{\mp}(b)$. Similarly, given a neighborhood $U^{\prime}$ of $\mathcal{F}^{+}(b)\cap\mathcal{F}^{-}(a)$ in the quadrilateral bounded by the leaves of $a,b$, for sufficiently large $m,n$ the map $\beta^{m}\alpha^{n}$ has a fixed point $z^{\prime}$ in $U^{\prime}$. Since $\beta^{m}\alpha^{n}=\beta^{m}(\alpha^{n}\beta^{m})\beta^{-m}$, we have that $z^{\prime}=\beta^{n}(z)$, as desired. ∎
@@
One easily shows that, up to shrinking the product neighborhood in the above corollary, the closed sets $C_{i}$ are either a closed interval, a Cantor set or an isolated point. The work in the next few sections will show that, as in the case of Anosov flows, only four of the six (up to symmetry) possible topological types of product can happen: One may have a product $\mathrm{Interval}\times\mathrm{Interval}$ (which happens only if the action is transitive), a product $\mathrm{Interval}\times\mathrm{Cantor}$ (which happens in extremal Smale classes, see Observation 6.6), a product $\mathrm{Cantor}\times\mathrm{Cantor}$ (which happens for any non-isolated point in a non-extremal Smale class), and finally isolated fixed points, i.e., of the form $\mathrm{Isolated}\times\mathrm{Isolated}$. See Proposition 4.12 and Corollary 4.14 for a proof.
@@
Using this, we will be able to prove Theorem 1.7. Recall the statement:
See 1.7
While the fact that $G$ acts topologically transitively on each singular Smale class is tautologically true, in order to deduce the “moreover…” part of the above result, one needs to deal separately with the case of Anosov-like actions such that all Smale classes are singular. We use the following result, which will also play a role in Proposition 6.3.
Lemma 3.15.
If $\overline{\mathrm{Fix}}_{G}$ consists only of isolated prong singularities, then these must lie in infinitely many distinct $G$-orbits. Consequently, there are infinitely many Smale classes.
We remark that there do exist examples of such Anosov-like actions with only isolated prong singularities as fixed points, see Proposition 8.10. To streamline the proof of this lemma, we use one more fact whose proof we postpone to Section 4. There is no circularity here since Lemma 3.15 and its consequence are not used in Section 4.
Lemma 3.16.
If $p$ is an isolated prong singularity, then each of its quadrants admits a lozenge with corner $p$.
This is proved in Proposition 4.12.
Proof of Lemma 3.15.
Suppose that $\overline{\mathrm{Fix}}_{G}$ consists only of isolated prong singularities.
By Lemma 3.16, each fixed point is a corner of a lozenge in each of its quadrants, thus each lozenge is included in a scalloped region in $P$. Fix such a lozenge $L_{1}$ with corner $a_{1}$, in scalloped region $S$ and enumerate the quadrants of $a_{1}$ by $Q^{1}_{1},Q^{1}_{2},\ldots Q^{1}_{k}$, so that $L_{1}\subset Q^{1}_{1}$. For concreteness, fix an orientation on this scalloped region $S$ so that $a_{1}$ is the upper right corner of $L$. Since $S$ can be realized in two ways as a line of lozenges, there exists $L_{2}\neq L_{1}$ in $S$ with $L_{1}\cap L_{2}\neq\emptyset$, and also with a corner $a_{2}$ in the upper right. (In fact, there are infinitely many such choices of $L_{2}$, we fix one). See Figure 4.
[…]
Figure 4. If $g_{1}(L_{1})=L_{2}$ then $g$ has a fixed point in $L_{1}\cap L_{2}$
If $a_{2}=g_{1}(a_{1})$ for some $g_{1}\in G$, and $g_{1}(L_{1})=L_{2}$ then we already have a contradiction: $g_{1}(L_{1})\cap L_{1}$ would contain a (necessarily nonsingular) fixed point for $g_{1}$. Thus, either $a_{2}$ is in a distinct $G$-orbit from $a_{1}$, or we have $a_{2}=g_{1}(a_{1})$ but $L_{2}$ is contained in $g_{1}(Q^{1}_{j})$ for some $j\neq 1$. We now iterate this process: $L_{2}$ is contained in a scalloped region $S_{2}$ and we may find a lozenge $L_{3}$ in $S_{2}$, with $L_{3}\cap L_{2}\cap L_{1}\neq\emptyset$, and with corner $a_{3}$ on the upper right. If we have $a_{3}=g_{2}(a_{1})$ (or $a_{3}=g_{2}(a_{2})$) for some $g_{2}\in G$, and $g_{2}(L_{1})=L_{3}$ (or $g_{2}(L_{2})=L_{3}$), then we arrive at the same contradiction.
Iterating this process, we get a sequence $a_{k}$ of singular points that are corners of lozenges $L_{n}$ such that $L_{1}\cap\dots\cap L_{n}\neq\emptyset$ for any $n$. If there are only finitely many distinct $G$-orbits of singular points, there is a subsequence $a_{k_{i}}$ that are all in the same $G$-orbit . Since $a_{k_{i}}$ has only finitely many quadrants, we will eventually find some $g=g_{k_{j}}g_{k_{i}}^{-1}$ which maps some $L_{k_{i}}$ to $L_{k_{j}}$, and so has a nonsingular fixed point, a contradiction. ∎
@@
Consider $\Lambda$ a regular Smale class. We will show that for any open sets $U,V$ such that $U\cap\Lambda\neq\emptyset$ and $V\cap\Lambda\neq\emptyset$, there exists $g\in G$ such that $gU\cap V\cap\Lambda\neq\emptyset$. This implies the existence of a dense orbit (the classical proof for continuous maps applies here, see e.g. [KH95, Lemma 1.4.2], since regular Smale classes are closed in the complement of the singular leaves in $P$).
@@
In the special case where there is a unique Smale class $\Lambda$, then $\Lambda$ is a regular Smale class by Lemma 3.15, so Axiom (A2) together with Proposition 3.9 implies that $\bar{\Lambda}=P$. Hence, $G$ is topologically transitive on $P$ and $P$ contains a dense subset of points fixed by nontrivial elements, i.e., Axioms (A2a) and (A2b) are both satisfied. Conversely, if $\mathrm{Fix}_{G}$ is dense, then $\mathcal{R}_{G}$ is dense (since $\mathcal{R}_{G}=\mathrm{Fix}_{G}\smallsetminus\mathrm{Sing}$, and $\mathrm{Sing}$ is a closed discrete subset of $P$ by Lemma 2.23), moreover any product neighborhood intersects at most one Smale class, so there is a unique Smale class by connectedness of $P\smallsetminus\mathrm{Sing}$. ∎
Theorem 1.5 is now an immediate consequence, since lifts of periodic orbits to $\widetilde{M}$ project to fixed points for the action of $\pi_{1}(M)$ on $\mathcal{O}_{\phi}$.
@@ 4. Smale chains and the structure of wandering sets
As described at the start of Section 3, for Anosov flows on compact manifolds, Smale classes correspond to basic sets. By Brunella [Bru93] these sets are separated by a collection of embedded tori transverse to the flow. Moreover, in the nontransitive case any basic set contains “boundary” orbits, which are periodic orbits such that (at least) one of their rays are contained in the wandering set (these are the free separatrices appearing in [BB02, Lemme 1.6]). See also [BL98] for a closely related theory for Smale diffeomorphisms of surfaces.
@@
Note that, by comparison, the first step of the standard proof of the existence of separating transverse tori is the existence of a Lyapunov function for Anosov flows. Furthermore, as mentioned in the introduction, in the case of pseudo-Anosov flows, here we detect more of the tori in the non-wandering set than what the proof using Lyapunov functions would, since Lyapunov functions only allow to separate chain transitive components, and not the components of the non-wandering set. See Propositions 8.8 and 8.9.
@@ 4.1. Structure of the regular and “wandering” sets
Equivalently, if $Q_{0},\dots,Q_{2p-1}$ are the $TL$ subquadrants, cyclically ordered, and $Q_{0}$ is wandering, then there exists $i$, $0\leq i\leq 2p-1$, such that both $Q_{i}$ and $Q_{i+1}$ are wandering.
@@
Up to renaming the $TL$ subquadrants, we assume that $Q_{0},\dots,Q_{2p-1}$ are ordered so that $Q_{0}$ shares a $\mathcal{F}^{+}$-side with $Q_{1}$ and a $\mathcal{F}^{-}$-side with $Q_{2p-1}$. If the $Q_{2i-1}$$i=1,\dots,p$ are all regular, then there exist points $y_{i}\in\mathrm{Fix}(G)\cap Q_{2i-1}$. As in the proof of Lemma 4.5, after applying a high positive or negative power of the element of $G$ fixing $c$ (which exists by axiom (A3)), we can replace $y_{i}$ with points $y_{i}^{\prime}$ with $\mathcal{F}^{\pm}{y_{i}^{\prime}}\cap\mathcal{F}^{\mp}(c)$ as close as we like to $c$. By then applying Corollary 3.11 using $y_{i}^{\prime}$ and $c$, we can obtain points $x_{i}\in\mathrm{Fix}(G)\cap Q_{2i-1}$ as close as we like to $c$. In particular, we can take these sufficiently close to $c$ so that $\mathcal{F}^{+}(x_{i})\cap\mathcal{F}^{-}(x_{i+1})\neq\emptyset$ and $\mathcal{F}^{+}(x_{p})\cap\mathcal{F}^{-}(x_{1})\neq\emptyset$. Let $h_{i}\neq id$ be an element of $G$ that fixes $x_{i}$. Then for appropriate powers $n_{i}$, we have that $h_{p-1}^{n_{p-1}}\dots h_{2}^{n_{2}}\mathcal{F}^{+}(x_{1})\cap\mathcal{F}^{-}(x_{p})\neq\emptyset$. That is, we are in the setting of Proposition 3.9, and therefore there is a point fixed by some nontrivial element of $G$ arbitrarily close to $\mathcal{F}^{+}(x_{p})\cap\mathcal{F}^{-}(x_{1})\in Q_{0}$, contradicting the fact that $Q_{0}$ is wandering. ∎
@@
Suppose $x\in\overline{\mathrm{Fix}}_{G}$ has a wandering TL-substrip $S$ based on a ray $r_{0}$ of a leaf of $\mathcal{F}^{+}$. Then $r_{0}$ makes a perfect fit with a leaf of $\mathcal{F}^{-}$ on the side of the wandering substrip.
@@
By Lemma 4.8, we may assume that the boundary of $S$ consists of $r_{0}$, a ray $r^{\prime}_{1}$ of a leaf fixed by some element $g\in G$, a compact interval of a leaf of $\mathcal{F}^{-}$ between the two initial points of the rays, and a part of a leaf $l^{-}\in\mathcal{F}^{-}$. Moreover, Lemma 4.8 gives three possibilities for $l^{-}$. One of the three options is that $l^{-}$ makes a perfect fit with $r^{+}$, which is what we want to show. Hence, we have to eliminate the other two.
@@ 4.2. Boundary leaves
A boundary leaf is a leaf $l\in\mathcal{F}^{\pm}$ such that there exists $a\in l\cap\overline{\mathrm{Fix}}_{G}$ and $a$ is a boundary point of $\overline{\mathrm{Fix}}_{G}\cap\mathcal{F}^{\mp}(a)$. Analogously, one can also define boundary faces for singular leaves.
@@
Let $x\in\mathcal{R}_{G}$ be such that $\mathcal{F}^{\pm}(x)$ is an isolated boundary leaf, i.e., for some small enough product neighborhood $U$ of $x$, we have $\overline{\mathrm{Fix}}_{G}\cap U\subset\mathcal{F}^{\pm}(x)$. Then $x\in\mathrm{Fix}_{G}$ and it is the corner of four wandering lozenges. In particular, the Smale class of $x$ consists only of $x$ and its images under $G$.
Similarly, if $p$ is an isolated prong singularity, then all quadrants of $p$ are wandering lozenges.
@@
Let $U$ be a product neighborhood of $x$ such that $\overline{\mathrm{Fix}}_{G}\cap U\subset\mathcal{F}^{+}(x)$. (The proof for $\mathcal{F}^{-}(x)$ being isolated is obviously symmetric). Since $x\in\mathcal{R}_{G}$, it is accumulated by elements of $\mathrm{Fix}_{G}$, but any such element in $U$ must be in $\mathcal{F}^{+}(x)$ by assumption. Since any leaf contains at most one element of $\mathrm{Fix}_{G}$ (see Remark 1.6), we deduce that $x\in\mathrm{Fix}_{G}$. Hence $\overline{\mathrm{Fix}}_{G}\cap U=\{x\}$, so (by the product structure of $\overline{\mathrm{Fix}}_{G}$) every TL subquadrant of $x$ are wandering and the conclusion follows from Corollary 4.10.
Similarly, if $p$ is an isolated prong singularity, Corollary 3.11 implies that each of its TL subquadrant must be wandering and the conclusion follows again from Corollary 4.10. ∎
@@
Assuming $y=\infty$, by the above, both TL subquadrants of $x$ along the ray $(x,y)$ are wandering. In particular, the ray $(x,y)$ is the base of wandering TL-substrips on each sides. By Lemma 4.9 there exist leaves $l_{1},l_{2}$ of $\mathcal{F}^{-}$ making perfect fits on each side. These are nonseparated in the leaf space of $\mathcal{F}^{+}$, so are fixed by some nontrivial element $g\in G$ (Prop. 2.12), forming the sides of adjacent $g$-invariant lozenges (Corollary 2.21), with common shared side on $\mathcal{F}^{+}(x)$. Moreover, these lozenges are wandering, so the common corner on $\mathcal{F}^{+}(x)$ is the element in $\overline{\mathrm{Fix}}_{G}\cap\mathcal{F}^{+}(x)$ closest to the perfect fit, i.e., it is $x$. ∎
@@
    II is unbounded, $a$ is fixed by $g$, and $a$ is the corner of two wandering lozenges with $I$ as their shared sides,